# fold/.style=folder, grow’=0, s sep=0.1pt, arr/.style=edge=-latex

### v2

@@ II-A Quantum Unitary Synthesis
To quantify the optimality of T-gate usage, we adopt the T-count metric, $\mathcal{T}(U)$. It is defined as the minimum number of T-gates required to implement a unitary $U$ up to a global phase for unitaries that can be exactly synthesized from the Clifford+T gate set [25, 26]. Finding the T-count, however, for multi-qubit circuits is still an open problem.
@@ III-A Classical Unitary Synthesis
The first heuristic method aimed at minimizing the T-gate count in synthesis using Clifford+T was introduced by Gosset et al. [25]. This method leveraged the channel representation of unitaries and a decision algorithm for determining the T-count of a unitary $U$, denoted as $\mathcal{T}(U)\leq m$, for a specified integer $m$. The complexity of this algorithm was characterized as $\mathcal{O}(2^{nm}\text{poly}(m,2^{n}))$, with $n$ representing the number of qubits and $m$ the T-count.
Building upon the framework established in Ref. [10], Mosca and Mukhopadhyay developed MIN-T-SYNTH, an algorithm that determines the T-count of a unitary by innovatively pruning the search space [11]. Despite its advancements, the algorithm’s exponential complexity relative to the T-count remains a challenge.

### v3

@@ II-A Quantum Unitary Synthesis
To quantify the optimality of T-gate usage, we adopt the T-count metric, $\mathcal{T}(U)$. It is defined as the minimum number of T-gates required to implement a unitary $U$ up to a global phase for unitaries that can be exactly synthesized from the Clifford+T gate set [25, 26, 27].
@@ III-A Classical Unitary Synthesis
The first deterministic method aimed at finding the T-count of a unitary which is exactly synthesizable using Clifford+T gate set was introduced by Gosset et al. [25]. This method leveraged the channel representation of unitaries and a decision algorithm for determining the T-count of a unitary $U$, denoted as $\mathcal{T}(U)\leq m$, for a specified integer $m$. The complexity of this algorithm was characterized as $\mathcal{O}(2^{nm}\text{poly}(m,2^{n}))$, with $n$ representing the number of qubits and $m$ the T-count.
The first heuristic algoroithm for synthesizing a T-count optimal unitary was developed by Mosca and Mukhopadhyay [11]. Their algorithm, MIN-T-SYNTH, prunes the search space by exploiting the structure of the channel representation of the unitary, achieving a conjectured polynomial complexity relative to the T-count. Despite its advancements, the algorithm’s exponential complexity relative to the number of qubits remains a challenge.
@@
Finally, the work in Ref. [44] used decision diagram representation of a unitary matrix [45] to synthesize exactly representable unitaries with the Clifford+T gate set. Their algorithm was able to synthesize unitaries with up to 15 qubits and millions of T-gates.